# Trial Length, Pricing, and Rationally Inattentive Customers

### v2

@@ 3.6 Aggregate Outcomes and Firm Objective
[…]
Since the cancellation probability $q^{*}$ is constant for all consumers with $v<P$, this expression simplifies to $\text{IR}(T,P)=P\cdot F(P)\cdot[1-q^{*}(P,\lambda(T))]$. This component of profit is central to our analysis, as it is the channel through which the firm’s choice of $T$ affects its revenue.
@@
Thus, we define the firm’s total profit is the sum of standard revenue from willing subscribers, as in standard economic models, and inattentive revenue from (rationally) “inattentive subscribers”.
@@
Equation 4 defines the ex-ante expected utility for a consumer considering the contract, averaged over the entire population. It is a weighted sum of the outcomes for two distinct types of consumers.
[…]
This comprehensive utility function means that consumers in our model are sophisticated and forward-looking. When deciding whether to sign up (Stage 1), they do not just consider the price $P$; they also anticipate the future hassle. They understand that a long trial (high $T$) or a high price (high $P$) will increase both their risk of a costly mistake and the cognitive burden of avoiding it. For consumers to be willing to sign up for the trial, this aggregate expected utility must be non-negative. This gives rise to an individual rationality (IR) or participation constraint, $U(T,P)\geq 0$, which is the central constraint on the firm’s behavior.
@@
We define the IR-Slack the marginal harm to consumer utility from extending the trial. It represents the "cost" to the firm, in terms of reduced offer attractiveness, of making the contract more cognitively burdensome. Applying the envelope theorem to (4), the only term that changes with $T$ through $\lambda(T)$ is the cognitive cost term.
[…]
The IR-Slack is the critical countervailing force that disciplines the firm’s incentive to extend trials indefinitely. The consumers in our model are sophisticated; they are not passive victims of their own inattention but are forward-looking agents who anticipate the cognitive costs and potential for error associated with managing a long-term commitment. When a firm increases the trial length $T$, consumers recognize that this makes the contract more difficult and costly to manage. This anticipated "hassle cost" reduces their ex-ante valuation of the firm’s offer.
The IR-Slack precisely measures this marginal erosion of consumer utility. A firm that ignores this effect will find its offer becomes unattractive, leading to lower sign-up rates and a smaller customer base. This finding aligns perfectly with the large-scale field experiment by Miller et al. (2023), who find that auto-renewal offers (which have a higher perceived hassle cost) see significantly lower take-up rates than auto-cancel offers. The IR-Slack therefore represents the cost side of the firm’s trade-off, a cost paid in the currency of consumer goodwill and participation. The formula is intuitive: the marginal harm to consumers is proportional to the rate of attention decay ($\beta/\lambda_{0}$), the cognitive effort required to overcome it (captured by $-H(q^{*})$), and the size of the population that is vulnerable to this friction ($F(P)$).
@@
$\max_{T,P}\;\;\Pi(T,P)\quad\text{such that}\quad U(T,P)=0.$
@@
4 Functional-Form-Free Results
In this section, we derive the foundational results of our model. These propositions are "functional-form-free" in the sense that they do not depend on the specific Shannon entropy cost function, but rather on a minimal set of qualitative properties that any reasonable model of costly attention should exhibit. Specifically, we require only that (a) the consumer’s ability to pay attention, $\lambda(T)$, decreases as the task’s deadline, $T$, moves further into the future, and (b) the consumer’s chosen probability of success, $q^{*}$, increases with their ability to pay attention, $\lambda$. These general results reveal the fundamental economic tensions that drive the firm’s strategic decisions and provide a robust theoretical backbone for the sharper, more specialized predictions that will follow.
@@ 4 Functional-Form-Free Results
A longer trial period unambiguously decreases the consumer’s effort to cancel. For any given renewal price $P>0$, if the rate of attention decay is positive ($\beta>0$), then an increase in trial length $T$ leads to a strict decrease in the optimal cancellation probability $q^{*}(P,\lambda(T))$.
[…]
@@
The proof is a direct application of the chain rule. The change in the optimal cancellation probability with respect to the trial length is given by $\frac{\partial q^{*}}{\partial T}=\frac{\partial q^{*}}{\partial\lambda}\frac{d\lambda}{dT}$. From the consumer’s optimization problem in Lemma 1, we established that a higher attention sensitivity $\lambda$ makes it cheaper to exert mental effort, leading to a higher chosen success probability, so $\frac{\partial q^{*}}{\partial\lambda}>0$. From our core behavioral assumption on attention decay, LABEL:ass:MA3, a longer trial reduces attention sensitivity, so $\frac{d\lambda}{dT}<0$ for any $\beta>0$. The product of a positive term and a negative term is necessarily negative, completing the proof. ∎
@@
This proposition establishes the primary behavioral channel through which the firm’s contract design influences consumer outcomes. The trial length $T$ is not merely a passive waiting period; it is an active strategic instrument that shapes the consumer’s cognitive environment. By extending the trial, the firm makes the future cancellation task less salient and more susceptible to being forgotten. This effectively raises the cognitive price of remembering. The rationally inattentive consumer, who is aware of this increased difficulty, optimally responds by reducing their monitoring effort, which manifests as a lower probability of successful cancellation. This result provides the firm with a direct and predictable lever to influence the rate of what we term "accidental renewals.". This analytical result also fits well with industry reports, such as RevenueCat, a top revenue management platform for subscription apps [^3] State of Subscription Apps 2025., as shown in Figure 3.
@@
By definition, inattentive revenue is $\text{IR}(T,P)=P\cdot F(P)\cdot[1-q^{*}(P,\lambda(T))]$. Holding the price $P$ constant, the only term in this expression that depends on the trial length $T$ is the cancellation probability $q^{*}$. From Proposition 1, we know that $\partial q^{*}/\partial T<0$. This implies that the probability of failing to cancel, which is $1-q^{*}$, must be strictly increasing in $T$. Since $P>0$ and we assume $F(P)>0$, the entire expression for IR must be strictly increasing in $T$. ∎
This result formalizes the firm’s primary incentive to offer long trial periods, an incentive entirely divorced from consumer learning. From a purely revenue-focused perspective, a longer trial is always better. It systematically increases the number of consumers who, despite preferring to cancel, fail to do so and are subsequently charged the renewal price. This finding provides a clear, attention-based explanation for a key stylized fact: firms often offer trial periods (e.g., 30 or 60 days) that far exceed the time necessary to evaluate the product’s features. If this were the only force at play in our model, the firm’s optimal strategy would be to set an infinitely long trial period to maximize the effects of consumer forgetfulness. The existence of finite trial periods in reality points to a crucial countervailing force.
Definition 4 (IR-Slack, revisited).
[…]
@@
If the attention decay rate is positive ($\beta>0$) and there is a positive mass of consumers who wish to cancel (per Assumption 2), there exists a finite optimal trial length $T^{*}\in(0,\infty)$that solves the firm’s optimization problem. This $T^{*}$ satisfies the first-order condition:
$\frac{\partial\Pi}{\partial T}=P\frac{\partial\text{IR}}{\partial T}-\text{IR-Slack}(T,P)=0.$
This is a central result of the paper. It establishes that the tension between exploiting consumer inattention and maintaining the attractiveness of the contract leads to a well-defined, finite optimal trial length. The firm’s decision rule is to increase the trial length as long as the marginal benefit from higher inattentive revenue ($\partial\text{IR}/\partial T$) exceeds the marginal cost from the erosion of consumer utility (IR-Slack). At $T=0$, the marginal benefit is typically high. As $T$ becomes very large, attention sensitivity $\lambda(T)$ approaches zero, consumers anticipate a near-certain failure to cancel, and the IR-Slack becomes prohibitively large. The optimal $T^{*}$ is found where these two marginal effects are perfectly balanced. This provides a clear, micro-founded rationale for the intermediate trial lengths commonly observed in the market, a phenomenon that is difficult to explain solely through learning models. This prediction also aligns with recent field evidence from Yoganarasimhan et al. (2023), who find that for a SaaS product, shortening a trial from 30 days to 7 days actually increased the subscription rate, suggesting the 30-day trial was already beyond the optimal length, consistent with our model’s prediction of an interior optimum.
@@
Consider an exogenous shock that makes attention uniformly cheaper, modeled as a shift in the attention sensitivity function from $\lambda(T)$ to $\lambda^{\prime}(T)=\gamma\lambda(T)$ with $\gamma>1$. For any given contract $(T,P)$:
@@
This proposition can be interpreted as analyzing the static effects of a pro-consumer policy like a "click-to-cancel" law or a technological innovation like a popular subscription management app. Such a change represents a clear welfare gain for consumers at a fixed contract. It lowers their cognitive costs of managing subscriptions and reduces the likelihood of costly mistakes, thereby increasing their expected utility from any given offer. For the firm, however, the immediate effect is negative on the inattention margin. The policy or technology directly undermines the firm’s ability to profit from consumer forgetfulness. This sets up a natural conflict of interest: firms may have an incentive to lobby against such regulations or design their platforms to be incompatible with third-party management tools, while consumer advocacy groups would champion them. This proposition provides the static foundation for the full equilibrium analysis of policy effects in Section 5.
@@
Fix a contract $(T,P)$ and let the population’s baseline attention sensitivity $\lambda$ be heterogeneous with distribution $G$ on $(0,\infty)$. Define the aggregate monetary loss from inattentive renewals as
[…]
where $q^{*}(P,\lambda)=1/(1+\exp(-\lambda P))$.
Suppose $G_{1}$ and $G_{2}$ have the same mean of $1/\lambda$ but $G_{2}$ is a mean-preserving spread of $G_{1}$ (i.e. $G_{2}$ places more probability mass in the extreme low- and high-attention tails). Then
@@
This proposition yields powerful implications for welfare analysis and policy design by examining the societal cost of “attention inequality.” The core of the result is that the individual probability of making a costly mistake, $1-q^{*}$, is a strictly convex function of the underlying cost of attention, $1/\lambda$. This mathematical property has a clear and compelling economic interpretation: interventions that improve attention have sharply diminishing returns.
Consider two types of consumers. The first is a "very attentive" consumer with a high $\lambda$ (and thus a low attention cost). They already have a high probability of remembering to cancel, $q^{*}$, so their failure probability, $1-q^{*}$, is close to zero. Making them even more attentive (e.g., by giving them a better calendar app) provides a negligible benefit; it is hard to improve upon near-perfection. The second is a "very inattentive" consumer with a low $\lambda$ (and a high attention cost). Their probability of failure, $1-q^{*}$, is substantial. A small improvement in their attention sensitivity can lead to a large reduction in their probability of making a costly mistake.
This convexity has three main implications. First, it means that from a social welfare perspective, policies or product designs that target the "worst-off" consumers in the attention space are far more cost-effective than uniform interventions. For example, a policy that provides simple subscription management tools to low-income or elderly populations (who may have lower average $\lambda$) could deliver outsized welfare gains compared to a broad public awareness campaign. Second, it suggests that markets with high heterogeneity in consumer attention are more vulnerable to exploitation. A market composed of a mix of extremely attentive and extremely inattentive individuals will suffer a greater aggregate monetary loss from this friction than a market where everyone has the same, average level of attentiveness. The gains from exploiting the highly inattentive more than offset the losses from failing to exploit the highly attentive. Finally, this provides a new lens through which to view market segmentation. A firm might find it profitable to identify and specifically target low-attention consumers with contracts that have particularly long trial periods or high renewal prices, as these are the consumers from whom the most inattentive revenue can be extracted. This raises important questions about fairness and the ethics of using data to target consumers based on their cognitive traits.

### v3

@@ 3.6 Aggregate Outcomes and Firm Objective
$\boxed{\;\text{IR}(T,P)=P\!\int_{0}^{P}\!\bigl[\,1-q^{*}(P,\tau(T))\bigr]f(v)\,dv\;}$ (2)
Since the cancellation probability $q^{*}$ is constant for all consumers with $v<P$, this expression simplifies to $\text{IR}(T,P)=P\cdot F(P)\cdot[1-q^{*}(P,\tau(T))]$. This component of profit is central to our analysis, as it is the channel through which the firm’s choice of $T$ affects its revenue.
@@
Hence, we define the firm’s total profit as the sum of standard revenue from willing subscribers, as in standard economic models, and inattentive revenue from (rationally) “inattentive subscribers”.
@@
The first term, $\int_{P}^{1}(v-P)f(v)dv$, represents the standard consumer surplus enjoyed by "happy subscribers." These are consumers whose valuation $v$ is greater than or equal to the price $P$. They intend to use the service post-trial and derive a net benefit from doing so. The next two terms capture the expected outcome for the "unhappy subscribers" (or more accurately, the "unwilling subscribers"), those with $v<P$ who would prefer to cancel. This term is itself composed of two distinct sources of disutility. The first component, $P\,F(P)\,\bigl[1-q^{*}(P,\tau(T))\bigr]$, is the expected monetary loss. It is the price $P$ multiplied by the probability of failing to cancel, aggregated over all consumers in this segment. The second component, $\frac{1}{\tau(T)}\,I\!\bigl(q^{*}(P,\tau(T))\bigr)\,F(P)$, represents the aggregate cognitive cost of attention. This is a crucial feature of our model: even consumers who successfully cancel (and thus pay nothing) still incur a utility loss from the mental effort required to ensure they remembered to do so.
This comprehensive utility function means that consumers in our model are sophisticated and forward-looking. When deciding whether to sign up (Stage 1), they do not just consider the price $P$; they also anticipate the future hassles. The sophisticated consumers understand that a long trial (high $T$) or a high price (high $P$) will increase both their risk of a costly mistake and the cognitive burden of avoiding it. For consumers to be willing to sign up for the trial, this aggregate expected utility must be non-negative. This gives rise to an individual rationality (IR) or participation constraint, $U(T,P)\geq 0$, which is the central constraint on the firm’s behavior.
@@
We define the IR-Slack as the marginal harm to consumer utility from extending the trial period. To derive this, we apply the envelope theorem to the consumer’s cost-minimization problem in Equation (3.5). The derivative of the minimized loss $L^{*}=(1-q^{*})P+C(q^{*};\tau)$ with respect to $\tau$ is:
$\frac{\partial L^{*}}{\partial\tau}=\frac{\partial}{\partial\tau}\left(\frac{1}{\tau}\mathcal{I}(q^{*}\parallel 1/2)\right)=-\frac{1}{\tau^{2}}\mathcal{I}(q^{*}\parallel 1/2).$
The derivative of aggregate utility with respect to $T$ is then:
$\frac{\partial U}{\partial T}=\frac{\partial U}{\partial\tau}\frac{d\tau}{dT}=\left(-F(P)\frac{\partial L^{*}}{\partial\tau}\right)\frac{d\tau}{dT}=\left(F(P)\frac{1}{\tau^{2}}\mathcal{I}(q^{*}\parallel 1/2)\right)\left(-\frac{\beta\tau^{2}}{\tau_{0}}\right)=-\frac{\beta}{\tau_{0}}F(P)\mathcal{I}(q^{*}\parallel 1/2).$
The IR-Slack is the negative of this quantity:
$\text{IR-Slack}(T,P):=-\frac{\partial U(T,P)}{\partial T}=\frac{\beta}{\tau_{0}}F(P)\,\mathcal{I}(q^{*}\parallel 1/2)>0.$
The above is a critical counterbalancing force that constrains the firm’s incentive to extend trials too long. The consumers in our model are sophisticated; they are not passive victims of their own inattention but are forward-looking agents who anticipate the cognitive costs and potential for error associated with managing a long-term commitment. When a firm increases the trial length $T$, consumers recognize that this makes the contract more difficult and costly to manage. This anticipated "hassle cost" reduces their ex-ante valuation of the firm’s offer.
The IR-Slack measures this marginal erosion of consumer utility. A firm that ignores this effect will find its offer becomes unattractive, leading to lower sign-up rates and a smaller customer base. This finding aligns perfectly with the large-scale field experiment by Miller et al. (2023), who find that auto-renewal offers (which have a higher perceived hassle cost) see significantly lower take-up rates than auto-cancel offers. The IR-Slack, therefore, represents the cost side of the firm’s trade-off, a cost paid in the currency of consumer goodwill and participation. The formula is intuitive: the marginal harm to consumers is proportional to the rate of attention decay ($\beta/\tau_{0}$), the cognitive effort required to overcome it (captured by $I(q^{*})$), and the size of the population that is vulnerable to this friction ($F(P)$).
@@
$\max_{T,P}\;\;\Pi(T,P)\quad\text{such that}\quad U(T,P)\geq 0.$
Because $U(T,P)$ is decreasing in both $T$ and $P$, the constraint binds at the optimum. Letting $\mu>0$ denote the Lagrange multiplier, the Lagrangian is $\mathcal{L}=\Pi(T,P)-\mu\,U(T,P)$.
@@
4 Functional Form Free Results
In this section, we derive the foundational results of our model. These propositions are "functional-form-free" in the sense that they do not depend on the specific Shannon entropy cost function, but rather on a minimal set of qualitative properties that any reasonable model of costly attention should exhibit. Specifically, we require only that (a) the consumer’s ability to pay attention, $\tau(T)$, decreases as the task’s deadline, $T$, moves further into the future, and (b) the consumer’s chosen probability of success, $q^{*}$, increases with their ability to pay attention, $\tau$. These general results reveal the main economic tensions that drive the firm’s strategic decisions and provide a robust theoretical backbone for the sharper, more specialized predictions that will follow.
@@ 4 Functional Form Free Results
A longer trial period unambiguously decreases the consumer’s effort to cancel. For any given renewal price $P>0$, if the rate of attention decay is positive ($\beta>0$), then an increase in trial length $T$ leads to a strict decrease in the optimal cancellation probability $q^{*}(P,\tau(T))$.
$\frac{\partial q^{*}(P,\tau(T))}{\partial T}\;<\;0.$
@@
The proof is a direct application of the chain rule. The change in the optimal cancellation probability with respect to the trial length is given by $\frac{\partial q^{*}}{\partial T}=\frac{\partial q^{*}}{\partial\tau}\frac{d\tau}{dT}$. From the consumer optimization problem in Lemma˜1, a higher attention sensitivity $\tau$ makes it cheaper for them to exert mental effort, leading to a higher chosen success probability, so $\frac{\partial q^{*}}{\partial\tau}>0$. From our core behavioral assumption on attention decay, Item˜MA-3, a longer trial reduces attention sensitivity, so $\frac{d\tau}{dT}<0$ for any $\beta>0$. The product of a positive term and a negative term is necessarily negative, thus complete the proof. ∎
@@
This proposition describes the primary behavioral channel through which the firm’s strategic plan design influences consumer outcomes. The trial length $T$ is not just a passive waiting period, it is an active strategic instrument that shapes the consumer’s cognitive environment. By extending the trial, the firm makes the future cancellation task less salient and more susceptible to being forgotten. This effectively raises the cognitive price of remembering. The rationally inattentive consumer, who is aware of this increased difficulty, optimally responds by reducing their monitoring effort, which manifests as a lower probability of successful cancellation. This result provides the firm with a direct and predictable approach to influence the rate of what we term "accidental renewals.". This analytical result also fits well with industry reports, such as RevenueCat, a top revenue management platform for subscription apps [^7] State of Subscription Apps 2025., as shown in Figure 3.
@@
By definition, inattentive revenue is $\text{IR}(T,P)=P\cdot F(P)\cdot[1-q^{*}(P,\tau(T))]$. Holding the price $P$ constant, the only term in this expression that depends on the trial length $T$ is the cancellation probability $q^{*}$. From Proposition˜1, we know that $\partial q^{*}/\partial T<0$. This implies that the probability of failing to cancel, which is $1-q^{*}$, must be strictly increasing in $T$. Since $P>0$ and we assume $F(P)>0$, the entire expression for IR must be strictly increasing in $T$. ∎
Intuitively, this rationalizes the firm’s primary incentive to offer long trial periods, an incentive cannot be fully explained by consumer learning. From a purely revenue-focused perspective, a longer trial is always better. It systematically increases the number of consumers who, despite preferring to cancel, fail to do so and are subsequently charged the renewal price. This finding establishes an attention-based explanation for a key stylized fact: firms frequently offer trial periods (e.g., 30-60 days) that far exceed the time necessary to evaluate the product’s features. If inattentive revenue were the only consideration and consumers signed up regardless of the burden of the trial, the firm’s profit would rise monotonically as the trial got longer, and the firm would be incentivized to extend trial length until the chance of forgetting is 100%. In reality, a longer trial also makes the offer less attractive ex ante, captured by the IR-Slack term, so the firm eventually loses participants. Balancing these two forces yields a finite, moderate optimal trial length.
@@
If the attention decay rate is positive ($\beta>0$) and there is a positive mass of consumers who wish to cancel (per Assumption˜2), there exists a finite optimal trial length $T^{*}\in(0,\infty)$that solves the firm’s optimization problem. This $T^{*}$ satisfies the first-order condition:
$\frac{\partial\mathcal{L}}{\partial T}=\frac{\partial\Pi}{\partial T}-\mu\frac{\partial U}{\partial T}=0$
which simplifies to:
$\frac{\partial\text{IR}}{\partial T}\;=\;\mu\;\text{IR-Slack}(T,P),\qquad\mu>0.$ (T‑FOC)
Equation (T‑FOC) states that the marginal inattentive revenue generated by a longer trial must be the same as the Lagrange–weighted marginal tightening of the participation constraint. This is a central result of our model. It establishes that the tension between exploiting consumer inattention and maintaining the attractiveness of the contract leads to a well-defined, finite optimal trial length. The firm’s decision rule is to increase the trial length as long as the marginal benefit from higher inattentive revenue ($\partial\text{IR}/\partial T$) exceeds the marginal cost from the erosion of consumer utility (IR-Slack). At $T=0$, the marginal benefit is typically high. As $T$ becomes very large, attention sensitivity $\tau(T)$ approaches zero, consumers anticipate a near-certain failure to cancel, and the IR-Slack becomes prohibitively large. The optimal $T^{*}$ is found where these two marginal effects are perfectly balanced. This provides a clear, micro-founded rationale for the intermediate trial lengths commonly observed in the market, a phenomenon that is difficult to explain solely through learning models. This prediction also aligns with recent field evidence from Yoganarasimhan et al. (2023), who find that for a SaaS product, shortening a trial from 30 days to 7 days actually increased the subscription rate, suggesting the 30-day trial was already beyond the optimal length, consistent with our model’s prediction of an interior optimum.
@@
Consider an exogenous shock that makes attention uniformly cheaper, modeled as a shift in the attention sensitivity function from $\tau(T)$ to $\tau^{\prime}(T)=\gamma\tau(T)$ with $\gamma>1$. For any given contract $(T,P)$:
@@
This proposition can be interpreted as analyzing the static effects of a pro-consumer policy like a "click-to-cancel" law or a technological innovation like a popular subscription management app. Such a change represents a clear welfare gain for consumers at a fixed contract. It lowers their cognitive costs of managing subscriptions and reduces the likelihood of costly mistakes, thereby increasing their expected utility from any given offer. For the firm, however, the immediate effect is negative on the inattention margin. The policy or technology directly undermines the firm’s ability to profit from consumer forgetfulness. This sets up a natural conflict of interest: firms may have an incentive to lobby against such regulations or design their platforms to be incompatible with third-party management tools, while consumer advocacy groups would champion them. This proposition provides the static foundation for the full equilibrium analysis of policy effects in Section˜5.
@@
Fix a contract $(T,P)$ and let the population’s baseline attention sensitivity $\tau$ be heterogeneous with distribution $G$ on $(0,\infty)$. Define the aggregate monetary loss from inattentive renewals as
$\mathcal{L}(G)=P\int_{0}^{P}\int_{0}^{\infty}\bigl[1-q^{*}(P,\tau)\bigr]\,f(v)\,dG(\tau)\,dv=P\,F(P)\!\int_{0}^{\infty}\bigl[1-q^{*}(P,\tau)\bigr]\,dG(\tau),$
where $q^{*}(P,\tau)=1/(1+\exp(-\tau P))$.
Suppose $G_{1}$ and $G_{2}$ have the same mean of $1/\tau$ but $G_{2}$ is a mean-preserving spread of $G_{1}$ (i.e. $G_{2}$ places more probability mass in the extreme low- and high-attention tails). Then
@@
This proposition yields powerful implications for welfare analysis and policy design by examining the societal cost of “attention inequality.” The core of the result is that the individual probability of making a costly mistake, $1-q^{*}$, is a strictly convex function of the underlying cost of attention, $1/\tau$. This mathematical property has a clear and compelling economic interpretation: interventions that improve attention have sharply diminishing returns.
Consider two types of consumers. The first is a "very attentive" consumer with a high $\tau$ (and thus a low attention cost). They already have a high probability of remembering to cancel, $q^{*}$, so their failure probability, $1-q^{*}$, is close to zero. Making them even more attentive (e.g., by giving them a better calendar app) provides a negligible benefit; it is hard to improve upon near-perfection. The second is a "very inattentive" consumer with a low $\tau$ (and a high attention cost). Their probability of failure, $1-q^{*}$, is substantial. A small improvement in their attention sensitivity can lead to a large reduction in their probability of making a costly mistake.
This convexity has three main implications. First, it means that from a social welfare perspective, policies or product designs that target the "worst-off" consumers in the attention space are far more cost-effective than uniform interventions. For example, a policy that provides simple subscription management tools to low-income or elderly populations (who may have lower average $\tau$) could deliver outsized welfare gains compared to a broad public awareness campaign. Second, it suggests that markets with high heterogeneity in consumer attention are more vulnerable to exploitation. A market composed of a mix of extremely attentive and extremely inattentive individuals will suffer a greater aggregate monetary loss from this friction than a market where everyone has the same, average level of attentiveness. The gains from exploiting the highly inattentive more than offset the losses from failing to exploit the highly attentive. Finally, this provides a new lens through which to view market segmentation. A firm might find it profitable to identify and specifically target low-attention consumers with contracts that have particularly long trial periods or high renewal prices, as these are the consumers from whom the most inattentive revenue can be extracted. This raises important questions about fairness and the ethics of using data to target consumers based on their cognitive traits.